# Qualitatively altered driven Dicke superradiance in extended systems due to infinitesimal perturbations

### v1

@@ II.1 Equations of motion, perfect Driven Dicke Model
In this work, we use transmons coupled to a 1D waveguide to build up the extended system and model the light by classical continuous wave driving. Because the energy gap between the lowest two levels and the anharmonicity are much greater than the linewidth, the transmons are approximated by two-level systems[80]. Throughout the paper, we use |$\ket{0},\ket{1}$ to represent the ground state and first excited state of a transmon. The operators involved in this work are defined as below:
@@
II.2 Observation
By [3, 95], the photon emission rate consists of two parts:
$\gamma(t)=\gamma_{0}(t)+\gamma_{SR}(t)$ (14)
where the first term
$\gamma_{0}(t)=\sum_{n}\Gamma\langle\hat{e}_{n}\rangle(t)$ (15)
refers to the individual spontaneous emission and the second term
II.3 Perturbing the driven Dicke model
In general, there are three ways to perturb the driven Dicke model. One way is to introduce a dissipative (or $\Delta\mathcal{L}$ type) perturbation:
@@ II.3 Perturbing the driven Dicke model
In this work, we investigate one example of this type of perturbation, the local dephasing with homogeneous dephasing rate for each transmon [94]:
@@
We investigate two examples of this type of perturbation: introducing individual driving phases
@@ III Results
In this section, we first illustrate the driven Dicke superradiance in a 1D lattice, using homogeneous Rabi frequencies $\Omega_{n}=\Omega=8\Gamma$, the detunings $\Delta_{n}=\Delta=0$, and equal separation between adjacent transmons $d=1\lambda$. For each later subsection, we keep everything the same except for one perturbation type. For all the plots of time evolution, unless otherwise stated, we scale the time $t$ to the decay rate $\Gamma$ and define the scaled time $\tau$ as:
@@ III.1 Perfect driven Dicke model: superradiance in a 1D array
Here we normalize the correlated photon emission rate by the decay rate, $\Gamma$, and the number of transmons, $N$. All the curves have fast oscillation at early times from the collective Rabi oscillation, and reach steady states after about $t=10/\Gamma$. Note that the steady-state normalized correlated photon emission rate $\gamma_{SR}/(N\Gamma)$ is proportional to the qubit number, $N$. Such linear relation implies the quadratic relation between the total correlated photon emission rate $\gamma_{SR}$ and $N$, which is a signature of the driven Dicke superradiance.
@@ III.2 Perturbation: local dephasing
For a 1D-lattice consisting of 4 transmons, we plot the early-time behavior of the correlated photon emission rates in Fig. 2 for different dephasing rates in Eq. (18). At early times, the curves with different dephasing rates overlap because the chosen dephasing rates are small. They gradually split from each other later, so dephasing can be considered as a perturbation at early times because the size of the change in the signal is proportional to a power of the perturbation. Similarly, the perturbations discussed below (driving phase, separation, and detuning) hardly affect the early-time behavior of the correlated photon emission rate as well.
Figure 2: Early-time behavior of $\gamma_{SR}/(N\Gamma)$ with dephasing rates $\Gamma_{\phi}=0$, $\Gamma/1000$, $\Gamma/500$, $\Gamma/200$, and $\Gamma/100$
For the long-term behavior of the system, we plot the time evolution of each transmon’s normalized correlated photon emission rate under different dephasing rates $\Gamma_{\phi}$ in Fig. 3. We use the log scale for the y-axis. In addition, for each curve we further scale $\Gamma t$ by a factor proportional to the dephasing rate to obtain the scaled time $\tau^{\prime}$ in Fig. 3, illustrated in the label on the top right of the figure. The straight lines in the plot imply that the dephasing causes exponential decay of the correlated photon emission rate for this time scale. We find that even a tiny dephasing rate ($\Gamma_{\phi}=0.001\Gamma)$ will destroy the superradiance asymptotically. With greater perturbation, the correlated photon emission rate decays faster. By scaling the time linearly to the dephasing rate, the curves overlap with each other. This implies that the magnitude of the perturbation does not significantly change the steady state, instead, it modulates the rate to reach the steady state.
Figure 3: Evolution of $\gamma_{SR}/(N\Gamma)$ with different dephasing rates. For each curve, the scaled time $\tau^{\prime}=\xi\Gamma t$, where $\Gamma_{\phi}=\xi\Gamma/1000$.
@@ III.3 Perturbation: driving phase
Introducing a driving phase that depends on the transmon renders the system driven out of the symmetric subspace. Here we investigate the effect of a linear driving phase perturbation:
@@
where $n$ is the index of the transmon, $\phi_{n}$ is the phase introduced to the $n$th transmon, $k_{\phi}$ is the slope of the driving phase and is considered a measure of the perturbation. The time evolution of the correlated photon emission rate is shown in Fig. 4. For each curve, we scale the original scaled time $\Gamma t$ by the square of the perturbation, illustrated in the label on the top right of the figure.
Figure 4: Same as Fig. 3 except with different slopes of the driving phase. For each curve, the scaled time $\tau^{\prime}=\xi^{2}\Gamma t$, where the perturbation $k_{\phi}=\xi/400$.
@@ III.4 Perturbation: separation
For this perturbation, the transmons are equally separated with lattice constant $d$ but the separation is slightly different from a wavelength. The perfect driven Dicke case has $d=1\lambda$ while the perturbed lattice constants will be $d=1\lambda+\Delta d$, where $\Delta d$, the deviation from the integer multiple of the wavelength, is a measure of the perturbation. Note that the perturbation is a combination of $\Delta\mathcal{L}$ and $\Delta H$ type perturbation, as discussed in Sec. II.3. By the Taylor expansion of Eq. (24) to second order, $\Delta\Omega_{nm}$, related to the $\Delta H$ type perturbation, is linear in $\Delta d$; while $\Delta\Gamma_{nm}$, associated with the $\Delta\mathcal{L}$ type perturbation, is quadratic in $\Delta d$.
Figure 5 shows the time evolution of the correlated photon emission rate with perturbation in lattice constant $\Delta d=0,\lambda/400,\lambda/200,\lambda/100,\lambda/50$. For each curve, we further scale its scaled time $\Gamma t$ by a factor proportional to the square of the perturbation, as illustrated in the labels on the top right of the figure.
@@
Figure 5: Same as Fig. 3 except with different lattice constants. For each curve, the scaled time $\tau^{\prime}=\xi^{2}\Gamma t$, where the perturbation $\Delta d=\xi\lambda/400$.
@@ III.5 An exceptional case: symmetric, and equally spaced detuning
The inhomogeneous detuning also leads to coupling outside of the symmetric state subspace. Here we will show the effect of linear detuning:
@@
where $n=0,1,2,...N-1$. The slope $k_{\Delta}$ is considered as a measure of the perturbation. Figure 6 illustrates the evolution of correlated photon emission rates. For each curve, the original scaled time $\Gamma t$ is further scaled by a factor to the square of the corresponding slope $k_{\Delta}$.
@@
Figure 6: Same as Fig. 3 except with different slopes of the detuning. For each curve, the scaled time $\tau^{\prime}=\xi^{2}\Gamma t$, where the perturbation $k_{\Delta}=\xi 0.05$.
@@
For the asymptotic behavior, the perturbation in detuning results in a negative correlated photon emission rate. In other words, the system becomes subradiant at steady state. As for the early-time behavior, each curve, with time scaled quadratically to the perturbation measured by the slope $k_{\Delta}$, overlaps with each other but splits later. The different strength perturbations have different steady states, unlike the perturbations in the previous section. Such phenomenon has been explained by the theory in [68, 69, 70] - because we introduce a symmetric profile of detunings so that for any transmon with detuning $\Delta_{n}$, there will be a corresponding transmon $n^{\prime}$ with the detuning $\Delta_{n^{\prime}}=-\Delta_{n}$. These two transmons can pair up to compose a so-called Dimer[68, 69, 70], which exhibits no coupling to the light field when they reach the steady state. Also, the experiment in [71] reports the so-called collectively induced transparency(CIT), where for enough many inhomogeneous emitters in a cavity driven by an external field, the contribution of each emitter with a certain detuning $\Delta_{n}$ to the cavity field scales as $\propto\Delta_{n}^{-1}$ and can pair with the emitter with the same amplitude but opposite detuning so as to cancel each other. As a result, the negative correlated photon emission rate cancels the spontaneous emission rate so that the total photon emission in Eq. (14) is zero.
@@ IV Explanation, an angular momentum coupling scheme
In this section, we will give an explanation for the behavior noted in Sec. III.
@@ IV.1 Diagonalization of the Liouvillian
For a constant Liouvillian matrix $\hat{\mathscr{L}}$, there must be at least one eigenvalue $\mu_{n}=0$ with a corresponding density operator with non-zero trace, indicating the steady state. There are a large number of eigenstates with $\mu_{n}=0$ for the perfect driven Dicke model indicating a large subspace of possible steady states. It is the degenerate, steady state subspace that is the key to understanding the behavior noted in Sec. III.
For a 4-transmon system, out of $(2^{4})^{2}=256$ eigenvalues, we only plot the cluster of 14 eigenvalues with the smallest real parts. We chose to plot these because the perfect driven Dicke model with $N=4$ has $14$ eigenvalues with $\mu_{n}=0$. The other eigenvalues, with more negative real parts, decay very fast and are not closely related to the long-term behavior. The eigenvalues are illustrated in Figs. 7, 8 and, 9 for the perturbation of dephasing, driving phase, and separation, respectively.
Figure 7: The cluster of 14 eigenvalues of the Liouvillian with the least negative real parts for the dephasing perturbation
@@
Figure 8: Same as Fig. 7 except for the driving phase perturbation
Figure 9: Same as Fig. 7 except for the separation perturbation
@@
There are two main features in the plots. First, there is a qualitative difference between Fig. 7 and Figs. 8 and 9. The real parts of eigenvalues are linear in the perturbation for the dephasing case. Given that the real parts indicate the rate of the decay behavior, this accounts for the linear dependence of the decay rate of $\gamma_{SR}/(N\Gamma)$ on the dephasing rate $\Gamma_{\phi}$ in Fig. 3. Similarly, the quadratic dependence of real parts of eigenvalues on the perturbation for the driven phase and separation cases, Figs. 8 and 9, accounts for the quadratic relation between the decay rate of $\gamma_{SR}/(N\Gamma)$ and the perturbation in the two cases ($\Delta d$ and $k_{\phi}$). Second, there is only one eigenvalue $\mu_{n}=0$ for non-zero perturbation and, when the perturbation approaches 00, all the $14$ eigenvalues become degenerate $\mu_{0}=\mu_{1}=...=\mu_{13}=0$. Since a zero eigenvalue means the corresponding eigenvector does not evolve under the equation of motion in Eqs. (28) and (30), the degeneracy of the eigenvalues with $\mu_{n}=0$, $N_{ss}$, indicates the possible number of steady-state density operators when there is no perturbation. To investigate the dependence of the degeneracy on the transmon number, we diagonalize the Liouvillian for $N$ ranging from $2$ to $7$ and Table 1 illustrates the degeneracy $N_{ss}$ as a function of the transmon number $N$.
@@
The symmetric and equally spaced detuning case is the exception. The diagonalization of the full Liouvillian matrix with $N=4$ shows $4$ zero eigenvalues even after the perturbation, in contrast to the unique zero eigenvalue in the dephasing, driving phase, and separation perturbations. The projection of the initial state, where all transmons are in the ground state, on the $4$ eigenstates yields the dimer state and the projection on the other states decay away. Because dimers do not couple to the waveguide mode, the total photon emission rate in Eq. (14) is exactly 00. So the correlated photon emission rate has the same magnitude but opposite sign to the spontaneous emission rate, which is proportional to the average excitation $\sum_{n}\langle\hat{e}_{n}\rangle$, as shown in Eq. (15). Different detuning profiles lead to different average excitations and account for the different steady-state correlated photon emission rates in Fig. 6.

### v2

@@ II.1 Equations of motion, perfect driven Dicke Model
To investigate the effect of perturbations on the perfect driven Dicke Model, it is necessary to study the dynamics of the unperturbed system. In this work, we use transmons coupled to a 1D waveguide to build up the extended system and model the light by classical continuous wave driving. Because the energy gap between the lowest two levels and the anharmonicity are much greater than the linewidth, the transmons are approximated by two-level systems[80]. Throughout the paper, we use |$\ket{0},\ket{1}$ to represent the ground state and first excited state of a transmon. The operators involved in this work are defined as below:
@@
II.2 Perturbing the driven Dicke model
Given the dynamics of the perfect driven Dicke model, we want to investigate the perturbed dynamics of the system and compare with the unperturbed one. In general, there are three ways to perturb the driven Dicke model. One way is to introduce a dissipative (or $\Delta\mathcal{L}$ type) perturbation:
@@ II.2 Perturbing the driven Dicke model
This type of perturbation leads to purely real contribution to the Louivillian superoperator which, in lowest order, contributes to the decay behavior. In this work, we investigate one example of this type of perturbation, the local dephasing with homogeneous dephasing rate for each transmon [94]:
@@
This type of perturbation leads to purely imaginary contribution to the Louivillian superoperator which, in lowest order, contributes to oscillation and contributes to decay rate in second order. We investigate two examples of this type of perturbation: introducing individual driving phases
@@ III Results
In this section, to show the effect of different perturbations, we first illustrate the driven Dicke superradiance in a 1D lattice, using homogeneous Rabi frequencies $\Omega_{n}=\Omega$, detunings $\Delta_{n}=\Delta=0$, and equal separation between adjacent transmons $d=1\lambda$ in Sec. III.1. In each later subsection, we keep everything the same except for one type of perturbation to compare the result with the unperturbed system. To guarantee the system is in the superradiant phase, according to [56, 57], we keep the ratio between the drive and the collective decay to be a constant ($\beta=2\Omega/(N\Gamma)=4$, that is, $\Omega=2N\Gamma$) for all the plots. We simulate for at most $8$ transmons and take the case for $N=4$ as an example in several figures. For all four types of perturbations, superradiance is nearly unchanged at early times, but it gradually transitions to a distinct value asymptotically. The duration of the superradiant behavior increases as the size of the perturbation decreases, with the scaling of this duration depending on the type of perturbation. It is worth mentioning that the phenomena observed in the $N=4$ case also extend to systems with more transmons, up to the maximum number of transmons in our calculation, $N=8$.
For all the plots of time evolution, unless otherwise stated, we scale the time $t$ to the decay rate $\Gamma$ and define the scaled time $\tau$ as:
@@
For the observation, by [3, 95], the photon emission rate consists of two parts:
$\gamma(t)=\gamma_{0}(t)+\gamma_{SR}(t)$ (23)
where the first term
$\gamma_{0}(t)=\sum_{n}\Gamma\langle\hat{e}_{n}\rangle(t)$ (24)
refers to the individual spontaneous emission and the second term
@@ III.1 Perfect driven Dicke model: superradiance in a 1D array
Here we normalize the correlated photon emission rate by the decay rate, $\Gamma$, and the number of transmons, $N$. All the curves have fast oscillation at early times from the collective Rabi oscillation, and reach steady states after about $t=10/\Gamma$. Note that the steady-state normalized correlated photon emission rate $\gamma_{SR}/(N\Gamma)$ is proportional to the qubit number, $N$. Such linear relation implies the quadratic relation between the total correlated photon emission rate $\gamma_{SR}$ and $N$, which is a signature of the driven Dicke superradiance. For the simulation with greater $N$, ranging from $5$ to $30$, this relation still holds.
@@ III.2 Perturbation: local dephasing
To investigate the effect of small local dephasing, we compare the time evolution of the perturbed system with the perfect driven Dicke model for both early-time and long-term behavior. The local dephasing does not conserve total spin $S$, leading to evolution out of the $S=N/2$ subspace. For a 1D-lattice consisting of $4$ transmons, the early-time behavior of the correlated photon emission rates is shown in Fig. 2 for different dephasing rates in Eq. (15).
At early times, the curves with different dephasing rates overlap because the chosen dephasing rates are much smaller than $\Gamma$ and $\Omega$. They gradually split from each other later. This strength of dephasing can be considered as a perturbation at early times because the size of the change in the signal is proportional to the strength of the perturbation. Similarly, the perturbations discussed below (driving phase, separation, and detuning) hardly affect the early-time superradiant behavior of the correlated photon emission rate when their strength is small.
Figure 2: Early-time behavior of $\gamma_{SR}/(N\Gamma)$ with dephasing rates $\Gamma_{\phi}=0$, $\Gamma/1000$, $\Gamma/500$, $\Gamma/200$, and $\Gamma/100$. The scaled time, $\tau=\Gamma t$
For the long-term behavior of the system, we plot the time evolution of each transmon’s normalized correlated photon emission rate under different dephasing rates $\Gamma_{\phi}$ in Fig. 3. We use the log scale for the y-axis. In addition, for each curve we scale the time $t$ by the dephasing rate $\Gamma_{\phi}$, to obtain the scaled time $\tau^{\prime}$ in Fig. 3. The straight lines in the plot imply that the dephasing causes exponential decay of the correlated photon emission rate for this time scale. In other words, while the positive correlated photon emission rate can persist for a limited time, even a small dephasing rate ($\Gamma_{\phi}=0.001\Gamma$) will eventually undermine the superradiance. With greater perturbation, the correlated photon emission rate decays faster. By scaling the time linearly to the dephasing rate, the curves overlap with each other. This implies that the magnitude of the perturbation does not significantly change the steady state, instead, it modulates the rate to reach the steady state.
For more transmons, the phenomenon above also holds: the correlated photon emission rates exhibit exponential decay and, with the same scaled time as Fig. 3, the curves with different perturbation strengths collapse, which means the decay rate of $\gamma_{SR}/(N\Gamma)$ is proportional to the perturbation strength. We also fixed the dephasing rate $\Gamma_{\phi}=0.01\Gamma$, calculated the time evolution of $\gamma_{SR}/(N\Gamma)$ for $N=3\sim 8$, and evaluated the population living in the $S=N/2$ subspace. The starting point of each curve’s exponential decay is proportional to $N$, but the slope does not significantly depend on $N$ and the asymptotic value of $\gamma_{SR}/(N\Gamma)$ is approximately zero, regardless of $N$. The population in the $S=N/2$ subspace decreases with $N$, which means local dephasing opens a path going out of the $S=N/2$ subspace and this effect becomes more significant when $N$ increases.
Figure 3: Evolution of $\gamma_{SR}/(N\Gamma)$ with different dephasing rates. We define $\tau^{\prime}=1000\Gamma_{\phi}t$, except for $\Gamma_{\phi}=0$ where $\tau^{\prime}=\Gamma t$.
@@ III.3 Perturbation: driving phase
Introducing a driving phase that depends on the transmon renders the system driven out of the $S=N/2$ subspace. Here we investigate the effect of a linear driving phase perturbation:
@@
where $n$ is the index of the transmon, $\phi_{n}$ is the phase introduced to the $n$th transmon, $k_{\phi}$ is the slope of the driving phase and is considered a measure of the perturbation. The time evolution of the correlated photon emission rate is shown in Fig. 4. For each curve, we scale the original scaled time $\Gamma t$ by the square of the perturbation, as shown in the caption of the figure.
Figure 4: Same as Fig. 3 except with different slopes of the driving phase. Except for the unperturbed ($k_{\phi}=0$) case, where $\tau^{\prime}=\Gamma t$, we define $\tau^{\prime}=(400k_{\phi})^{2}\Gamma t$.
@@
For the same calculation as Fig. 4 but with more transmons, the curves with smaller perturbation strengths overlap very well, but gradually deviate from each other when the perturbation strength grows. This effect becomes more significant when the number of transmon increases. One reason lies in the linear driving phase we introduce: this phase deviation accumulates with the number of transmons and will become too large to be considered as a perturbation when $N$ increases. When the perturbation strength is fixed at $k_{\phi}=0.02$ but $N$ grows from $3$ to $8$, the starting point of $\gamma_{SR}/(N\Gamma)$ rises up, as implied by Fig. 1, but the initial slope of the decay increases with $N$ and the asymptotic value scales approximately as $\sim 1/N$. Meanwhile, the population in the $S=N/2$ declines with $N$, that is, the driving phase perturbation drives the system out of the $S=N/2$ subspace and this is increasingly significant with larger $N$.
@@ III.4 Perturbation: separation
An imperfect separation also opens a path to drive the system out of the $S=N/2$ subspace. To investigate the effect of the perturbation due to the separation, we compare the long-term behavior of the correlated photon emission rate of the perturbed system with the perfect one. For this perturbation, the transmons are equally separated with lattice constant $d$ but the separation is slightly different from a wavelength. The perfect driven Dicke case has $d=1\lambda$ while the perturbed lattice constants will be $d=1\lambda+\Delta d$, where $\Delta d$, the deviation from the integer multiple of the wavelength, is a measure of the perturbation. Note that the perturbation is a combination of $\Delta\mathcal{L}$ and $\Delta H$ type perturbation, as discussed in Sec. II.2. By the Taylor expansion of Eq. (21) to second order, $\Delta\Omega_{nm}$, related to the $\Delta H$ type perturbation, is linear in $\Delta d$; while $\Delta\Gamma_{nm}$, associated with the $\Delta\mathcal{L}$ type perturbation, is quadratic in $\Delta d$.
Figure 5 shows the time evolution of the correlated photon emission rate with perturbation in lattice constant $\Delta d=0,\lambda/400,\lambda/200,\lambda/100,\lambda/50$. For each curve, we further scale its scaled time $\Gamma t$ by a factor proportional to the square of the perturbation, as shown in the caption of the figure.
@@
Similar to the driving phase case, for the same calculation as Fig. 5 except for more transmons, the curves corresponding to minor perturbation collapse but those with stronger perturbation gradually deviate. This effect becomes increasingly significant with larger $N$ because the maximum change in separation accumulates with the number of transmons and will become too large to be considered as a perturbation. When the perturbation strength is fixed at $\Delta d=0.02\lambda$ and $N$ varies from $3$ to $8$, $\gamma_{SR}/(N\Gamma)$ starts declining at a value proportional to $N$ but with a steeper slope and settles at a lower value, which scales as approximately $\sim 1/N$. At the same time, the population in the $S=N/2$ also goes down with higher $N$. This means the perturbation in separation couples the system outside of the $S=N/2$ subspace and this effect is enhanced with more transmons.
Figure 5: Same as Fig. 3 except with different lattice constants. Except for the unperturbed ($\Delta d=0$) case, where $\tau^{\prime}=\Gamma t$, we define $\tau^{\prime}=(400\Delta d/\lambda)^{2}\Gamma t$.
@@ III.5 An exceptional case: symmetric, and equally spaced detuning
Inhomogeneous detuning also leads to coupling outside of the $S=N/2$ subspace. When the detuning has symmetric changes, the perturbation is shown to be an exceptional case due to the special configuration of detuning profiles. Here we will show the effect of linear detuning:
@@
where $n=0,1,2,...N-1$. The slope $k_{\Delta}$ is considered as a measure of the perturbation. In contrast to the above $3$ perturbations, the long-term behavior of the correlated photon emission rate is negative, implying that the system becomes subradiant. To explicitly show the evolution of the extent of subradiance, we plot the total photon emission rate $\gamma/(N\Gamma)$ instead of the correlated photon emission rate in Fig. 6. The system is subradiant when this scaled total photon emission rate is less than $\simeq 1/2$. For each curve, the original scaled time $\Gamma t$ is further scaled by a factor to the square of the corresponding slope $k_{\Delta}$.
@@
Figure 6: Similar to Fig. 3 except with different slopes of the detuning and showing the scaled total photon emission rate. The inset shows the early-time behavior of $\gamma/(N\Gamma)$. Except for the unperturbed ($k_{\Delta}=0$) case, where $\tau^{\prime}=\Gamma t$, we define $\tau^{\prime}=(k_{\Delta}/0.05)^{2}\Gamma t$.
@@
Each curve shows a relative faster decay at early times, followed by an extremely slow tail. For the early-time behavior, each curve, with time scaled quadratically to the perturbation measured by the slope $k_{\Delta}$, overlaps with each other but splits later. For the asymptotic behavior, the total photon emission rate shows exponential decay. Note that even if the time is scaled quadratically to $k_{\Delta}$, the late-time exponential decay still grows with increasing $k_{\Delta}$. So this decay rate depends on the perturbation strength beyond the second order. Despite the extremely slow decay, we claim that the system becomes totally subradiant at steady state - the total photon emission rate $\gamma$ is approximately 00. This can be verified by evaluating the steady-state density operator using the null eigenvectors of the Liouvillian matrix and Eq. (30) discussed in Sec. IV.1. The zero total photon emission rate has been explained by the theory in [68, 69, 70] - because we introduce a symmetric profile of detunings so that for any transmon with detuning $\Delta_{n}$, there will be a corresponding transmon $n^{\prime}$ with the detuning $\Delta_{n^{\prime}}=-\Delta_{n}$. These two transmons can pair up to compose a so-called Dimer[68, 69, 70], which exhibits no coupling to the light field when they reach the steady state. Also, the experiment in [71] reports the so-called collectively induced transparency(CIT), where for enough many inhomogeneous emitters in a cavity driven by an external field, the contribution of each emitter with a certain detuning $\Delta_{n}$ to the cavity field scales as $\propto\Delta_{n}^{-1}$ and can pair with the emitter with the same amplitude but opposite detuning so as to cancel each other. As a result, the negative correlated photon emission rate cancels the spontaneous emission rate so that the total photon emission in Eq. (23) is zero. As for the steady-state correlated photon emission rates, we claim that they are different, unlike the perturbations in the previous sections. The reason lies in different detuning profiles induced by different $k_{\Delta}$, rendering different average excitations. By Eq. (24), the average spontaneous photon emission rates $\gamma_{0}/(N\Gamma)$ are equal to the average excitation and therefore are different. Because the total photon emission, which is the sum of the spontaneous photon emission rate and correlated photon emission rate, is zero, different spontaneous photon emission rate implies the different correlated photon emission rate.
With greater $N$, the decreasing of $\gamma_{SR}/(N\Gamma)$ starts at a higher value but with a sharper slope. The asymptotic behavior depends on whether $N$ is even or odd: when the transmon number is even, the tail becomes increasingly slow with larger $N$ and eventually becomes fully subradiant. In contrast, if the transmon number is odd, the system is not fully subradiant - the asymptotic total photon emission rate is not zero and decreases with $N$. The reason lies in the dangling transmon - the rest $N-1$ transmons pair up to generate dimers and are effectively decoupled from the field, leaving one transmon that can still emit photons. When $N$ increases, due to our specific perturbation profile, the dangling transmon is increasingly detuned from the driving, rendering less excitation and photon emission. We also fixed $k_{\Delta}=0.2$ and estimated the steady-state population in the $S=N/2$ subspace for $N=2\sim 6$. In contrast to the previous $3$ kinds of perturbations, the population is at least one order of magnitude smaller and shows alternative behavior - the population for the even $N$ case is not only lower but also decreases faster than the odd $N$ case. This coincides with the alternative asymptotic behavior of the total photon emission rate. However, whether $N$ is even or not, the population declines, which means the symmetric, linear detuning perturbation paves the way to leak out of the $S=N/2$ subspace and the effect is increasingly significant with more transmons.
@@ IV Explanation, an angular momentum coupling scheme
In this section, we will give an explanation for the behavior noted in Sec. III. We first diagonalize the Liouvillian matrix and capture the features of the eigenvalues, revealing a degenerate subspace of steady states. Then we investigate the degenerate subspace using an angular momentum basis.
@@ IV.1 Diagonalization of the Liouvillian
Considering that the Linbladian in the master equation is a superoperator, it is convenient to investigate the dynamics of the system if the equation of motion can be mathematically represented by a matrix-vector product form. This is particularly useful for time-independent Lindbladians like those studied here.
@@
For a constant Liouvillian matrix $\hat{\mathscr{L}}$, there must be at least one eigenvalue $\mu_{n}=0$ with the corresponding null eigenvector. By Eq. (30), when $t\rightarrow\infty$, all the eigenvectors with negative real parts decay to zero. So the linear combination of the null eigenvectors $\vec{\rho}_{n}$ yields the steady-state density operator. When there is more than one null eigenvector, the overlaps together with null eigenvectors determine the steady state. However, if there is only one null eigenvector, the steady state is solely determined by the null eigenvector, which is the only eigenvector with non-zero trace. Specifically, if the system is prepared so that its density operator can be reshaped into a null eigenvector, the system will stay in this state forever. The steady-state spontaneous photon emission rate and correlated photon emission rate can be evaluated by Eqs. (24) and (25), which will indicate whether the steady state is superradiant or subradiant. There are a large number of eigenstates with $\mu_{n}=0$ for the perfect driven Dicke model indicating a large subspace of possible steady states. Thus, the steady state for the perfect driven Dicke model depends on the initial state which we have taken to be the state with all transmons in their ground state. It is the degenerate, steady state subspace that is the key to understanding the behavior noted in Sec. III because the steady state depends on degenerate perturbation theory within the superoperator. For the systems in Secs. III B,C,D, a nonzero perturbation leads to a single null eigenvector so that the steady state is independent of the initial conditions.
For perturbations of the driven Dicke model, it is the perturbation theory within the degenerate null subspace that determines the long-time behavior. As is discussed in Sec. II.2, there are two types of perturbations: dissipative (or $\Delta\mathcal{L}$ type) perturbation and $\Delta H$ type perturbation, each contributing to purely real (by Eq. (14)) and purely imaginary perturbations (by Eq. (16)), respectively. For purely real perturbations, the eigenvalues will have a real part proportional to the perturbation strength. For purely imaginary perturbations, the eigenvalues will have imaginary part proportional to the perturbation strength and real part quadratic in the strength because the real part arises from second order in the perturbation.
For a 4-transmon system, out of $(2^{4})^{2}=256$ eigenvalues, we only plot the cluster of 14 eigenvalues with the smallest real parts. We chose to plot these because the perfect driven Dicke model with $N=4$ has $14$ eigenvalues with $\mu_{n}=0$. The other eigenvalues, with more negative real parts, decay quickly and are not closely related to the long-term behavior. The eigenvalues are illustrated in Figs. 7, 8 and, 9 for the perturbation of dephasing, driving phase, and separation, respectively.
Figure 7: The cluster of 14 eigenvalues of the Liouvillian with the least negative real parts for the dephasing perturbation.
@@
Figure 8: Same as Fig. 7 except for the driving phase perturbation.
Figure 9: Same as Fig. 7 except for the separation perturbation.
@@
There are two main features in the plots. First, there is a qualitative difference between Fig. 7 and Figs. 8 and 9. The real parts of eigenvalues are linear in the perturbation for the dephasing case. By Eq. (30), the negative real parts indicate the decay rate for the corresponding eigenvector and, in turn, the dynamics of the density operator. The linear dependence of the real parts on the perturbation strength means the decay rates contributing to the dynamics of the density operator are also linear to the perturbation strength. This accounts for the linear dependence of the decay rate of $\gamma_{SR}/(N\Gamma)$ on the dephasing rate $\Gamma_{\phi}$ in Fig. 3. Similarly, the quadratic dependence of real parts of eigenvalues on the perturbation for the driven phase and separation cases, Figs. 8 and 9, accounts for the quadratic relation between the decay rate of $\gamma_{SR}/(N\Gamma)$ and the perturbation in the two cases ($\Delta d$ and $k_{\phi}$). Second, there is only one eigenvalue $\mu_{n}=0$ for non-zero perturbation and, when the perturbation approaches 00, all the $14$ eigenvalues become degenerate $\mu_{0}=\mu_{1}=...=\mu_{13}=0$. Since a zero eigenvalue means the corresponding eigenvector does not evolve under the equation of motion in Eqs. (28) and (30), the degeneracy of the eigenvalues with $\mu_{n}=0$, $N_{ss}$, indicates the possible number of steady-state density operators when there is no perturbation. To investigate the dependence of the degeneracy on the transmon number, we diagonalize the Liouvillian for $N$ ranging from $2$ to $7$ and Table 1 illustrates the degeneracy $N_{ss}$ as a function of the transmon number $N$.
@@
Figure 10: Same as Fig. 7 except for the detuning perturbation. This range of $k_{\Delta}$ is beyond where lowest order perturbation is accurate to show the splitting of the four smallest eigenvalues.
The symmetric and equally spaced detuning case is the exception. The steady-state density operator, indicated by the unique null eigenvector, is a pure state where the transmons form dimers. Because dimers do not couple to the waveguide mode, the total photon emission rate in Eq. (23) is approximately 00. In Fig. 10, despite the uniqueness of the null eigenvalue, there are $3$ pure real eigenvalues ($\mu_{1}$, $\mu_{2}$, and $\mu_{3}$) very close to it - they are almost degenerate to the null eigenvalue until around $k_{\Delta}>0.13$. We find that these three eigenvalues are proportional to $k_{\Delta}^{4}$, indicating that the long-term decay in Fig. 6 arises from fourth order in perturbation theory. This explains the extremely slow tails for $k_{\Delta}=0.05$ and $k_{\Delta}=0.1$ compared to the relatively fast decay for $k_{\Delta}=0.2$ in Fig. 6. Different detuning profiles lead to different average excitations and account for the different steady-state correlated photon emission rates, as discussed below Fig. 6. Note that there are $4$ pairs of eigenvalues with the same real parts: $\mu_{2}$ and $\mu_{3}$, $\mu_{6}$ and $\mu_{7}$, $\mu_{8}$ and $\mu_{9}$, $\mu_{10}$ and $\mu_{11}$. Except for $\mu_{2}$ and $\mu_{3}$, which are real and degenerate, all the other $3$ pairs of eigenvalues are mutually conjugate.